solve and check thank you
Anna has 12 bills in her wallet, some $5 and some $10. The total value of the bills is $100. How many of each bill does Anna have?

Answer by ptaylor(1894)

Let x=number of $5 bills in Anna's wallet
Then 12-x=number of $10 bills in her wallet
Value of $5 bills=$5x
Value of $10 bills=$10(12-x)
Now we are told that the total value of the bills is $100. So----

$5x+$10(12-x)=$100 get rid of parens
$5x+$120-$10x=$100 subtract $120 from both sides
$5x+$120-$120-$10x=$100-$120 collect like terms
-$5x=-$20 divide both sides by -$5
x=4---------------------number of $5 bills
12-x=12-4=8----------------------------number of $10 bills
CK
$5(4)+$10(8)=$100
$20+$80=$100
$100=$100

Hope this helps---ptaylor